# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v2

@@ 3 The WellPlay Dataset
Script Agents Victims #token(CN) #token(EN) Question
@@
Table 1: Dataset Statistics. Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types. The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
WellPlay Dataset
We built an evaluation dataset, called WellPlay, from the existing Conan dataset (Zhao et al., 2024c), derived from background narratives created for MMGs. Conan labels relationships from each character’s perspective, along with ground truth relationships, consolidating all available information. To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers. Besides existing labels of interpersonal relationships that gauge agents’ understanding of the situation, we also recruited four annotators to label character objectives and reasoning. The dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims as detailed in Table 1 . Our evaluation questions are grounded in characters’ objectives in the game and are crafted with input from human players, focusing on the following question types:
(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
(2) Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
(3) Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
Evaluation Metrics
We report the performance of each type of question and also calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance. The awarded points follow the scoring system used in MMGs, which awards 10 points for objective questions, 5 points for reasoning questions, and 2 points for relation questions. For these calculations, we utilise weighted mean and weighted standard deviation methodologies. In addition to assessing the accuracy of the three types of questions and the overall score, we also evaluate the Win Rates of players for performance assessment. The game rules are as follows: every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) will be eliminated. Additionally, we assume that murderers would always vote for someone else instead of voting for themselves. If the true murderer is successfully voted out, the game is considered a victory for the players seeking to identify the culprit. The detailed about the dataset and calculations is provided in Appendix B.
@@ 4.1 Experimental Setup
Base Models
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 (Default) and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and variance. The detailed introduction to the experiment settings is in Appendix C.1.
Baselines
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games. Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games. We adapted their methodology to MMGs for a fair comparison, with implementation details in Appendix C.2.
(1) Werewolf (Xu et al., 2023). Werewolf is another multi-agent game, where players identify werewolves through group discussion. Questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. We adapt its pre-set game instructions and role-specific information to MMGs settings. (2) Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b). Agents think, reflect, and choose who and what to ask based on their objectives. We use the framework from previous works, only replacing the agents’ objectives with those set in MMGs. (3) ThinkThrice (Wu et al., 2024). Designed for MMGs, agents craft questions from retrieved memory and the current scenario. (4) Personal Perspective. For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script. (5) Omniscient Perspective. Agents do not actively participate but make their final decisions based on all agents’ scripts.
To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and PLAYER*). This allows us to fairly compare the performance of the side trying to find the murderer, who face the same level of opposition from the murderer’s side.
4.2 Results
Performance Evaluation Results
@@ 4.2 Results
Figure 5 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table A3). PLAYER* shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Notably, PLAYER* significantly outperforms others in objective questions, even surpassing OP, which has access to all information. The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies. It represents the ideal “search” endpoint. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model. This showcases the “approximate” ability of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective. Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
@@
The results demonstrate that PLAYER* outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability. The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals. This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
Our findings reveal a significant discrepancy between the win rate and the objective metric (the average of non-murderer players correctly identifying the murderer) for agents, with OP exhibiting a small difference and PP showing the largest difference. The case analysis unveils that in OP games, agents make similar decisions, resulting in voting outcomes that are either mostly correct or mostly wrong, leading to consistency between individual and collective voting results. Conversely, in PP games, agents base their decisions on their individual scripts, leading to independent actions and difficulty in reaching a consensus to vote out the murderer. The performance of the agents falls between that of OP and PP, with vote distributions more dispersed than OP but more concentrated than PP. O-CoT suffered from unfortunate circumstances, falling short by a single vote in numerous games. This resulted in O-CoT being on the verge of victory in many instances but ultimately yielding poor overall results. We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
Efficiency and Cost Comparison
@@
As shown in Figure 6, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table (Table A5). The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages. [^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Figure 6: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. Costs are reported for one complete gameplay and one evaluation process for each script.
@@
The costs are divided into Gameplay and Evaluation phases. Notably, PP and OP methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner. We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps. In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives. PLAYER* stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, PLAYER* minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase. In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset. Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in PP and OP costs arising from their differing levels of script access.

### v3

@@ 3 The WellPlay Dataset
MMG #Agents #Victims #token(CN) #token(EN) Question
@@
Table 1: Dataset Statistics. Agents is the count of players, Victims is the number of victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types. The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
Dataset
We built an evaluation dataset, WellPlay, derived from background narratives created for MMGs (Zhao et al., 2024c). To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers. We have recruited four annotators to label inferential questions on:
(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles the wallet, for each character in the game.
(2) Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, or disappearance); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, or manslaughter).
(3) Relations. This includes interpersonal relationships between victims and others, as well as relationships among suspects, with labels adapted from the Conan dataset (Zhao et al., 2024c).
WellPlay encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims (see Table 1). More details of the dataset and annotating schema are in Appendix B.
@@ 4 Experiments
Script Evaluation Werewolf O-CoT ThinkThrice Questum
Death Wears White (9 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$033_{\pm.047}.$$067_{\pm.047}.$$067_{\pm.047}.$$033_{\pm.047}.$
Reasoning .$258_{\pm.024}.$$324_{\pm.014}.$$327_{\pm.012}.$$356_{\pm.030}.$
Relations .$421_{\pm.024}.$$458_{\pm.041}.$$384_{\pm.026}.$$435_{\pm.017}.$
Ghost Revenge (7 players, 3 victims) Win Rate .$222_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$333_{\pm.000}.$
Objective .$193_{\pm.066}.$$158_{\pm.043}.$$193_{\pm.025}.$$333_{\pm.066}.$
Reasoning .$307_{\pm.008}.$$322_{\pm.019}.$$377_{\pm.022}.$$353_{\pm.016}.$
Relations .$314_{\pm.018}.$$295_{\pm.018}.$$353_{\pm.025}.$$353_{\pm.018}.$
Danshui Villa (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$083_{\pm.000}.$$111_{\pm.039}.$$194_{\pm.039}.$$111_{\pm.079}.$
Reasoning .$286_{\pm.027}.$$286_{\pm.019}.$$310_{\pm.004}.$$286_{\pm.004}.$
Relations .$312_{\pm.027}.$$365_{\pm.045}.$$259_{\pm.015}.$$296_{\pm.075}.$
Unfinished Love (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$500_{\pm.000}.$
Objective .$083_{\pm.118}.$$028_{\pm.039}.$$000_{\pm.000}.$$333_{\pm.068}.$
Reasoning .$443_{\pm.023}.$$426_{\pm.035}.$$481_{\pm.039}.$$536_{\pm.028}.$
Relations .$472_{\pm.034}.$$519_{\pm.036}.$$514_{\pm.020}.$$560_{\pm.007}.$
Cruise Incident (5 players, 1 victim) Win Rate .$667_{\pm.471}.$ $1.000_{\pm.000}$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective .$417_{\pm.118}.$$500_{\pm.000}.$$583_{\pm.312}.$$667_{\pm.236}.$
Reasoning .$458_{\pm.068}.$$444_{\pm.086}.$$458_{\pm.059}.$$528_{\pm.052}.$
Relations .$367_{\pm.082}.$$422_{\pm.042}.$$411_{\pm.016}.$$422_{\pm.016}.$
Sin (4 players, 1 victim) Win Rate .$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$
Objective .$333_{\pm.272}.$$000_{\pm.000}.$$000_{\pm.000}.$$444_{\pm.314}.$
Reasoning .$650_{\pm.108}.$$467_{\pm.024}.$$533_{\pm.047}.$$550_{\pm.041}.$
Relations .$333_{\pm.067}.$$571_{\pm.067}.$$413_{\pm.090}.$$492_{\pm.022}.$
Deadly Fountain (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Reasoning .$381_{\pm.000}.$$444_{\pm.098}.$$508_{\pm.022}.$$587_{\pm.022}.$
Relations .$250_{\pm.068}.$$194_{\pm.039}.$$222_{\pm.039}.$$333_{\pm.068}.$
Unbelievable Incident (5 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$083_{\pm.118}.$$000_{\pm.000}.$$000_{\pm.000}.$$083_{\pm.118}.$
Reasoning .$431_{\pm.052}.$$278_{\pm.071}.$$194_{\pm.071}.$$472_{\pm.020}.$
Relations .$622_{\pm.083}.$$733_{\pm.109}.$$289_{\pm.126}.$$533_{\pm.054}.$
Desperate Sunshine (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective .$333_{\pm.000}.$$111_{\pm.157}.$$556_{\pm.157}.$$778_{\pm.157}.$
Reasoning .$537_{\pm.026}.$$630_{\pm.094}.$$759_{\pm.052}.$$741_{\pm.069}.$
Relations .$491_{\pm.094}.$$574_{\pm.035}.$$491_{\pm.047}.$$556_{\pm.068}.$
Riverside Inn (4 players, 1 victim) Win Rate .$083_{\pm.118}.$$000_{\pm.000}.$$083_{\pm.118}.$$250_{\pm.204}.$
Objective .$117_{\pm.024}.$$117_{\pm.062}.$$117_{\pm.094}.$$383_{\pm.062}.$
Reasoning .$248_{\pm.013}.$$312_{\pm.022}.$$281_{\pm.011}.$$373_{\pm.017}.$
Relations .$531_{\pm.014}.$$541_{\pm.038}.$$483_{\pm.030}.$$589_{\pm.071}.$
Solitary Boat Firefly (6 players, 4 victims) Win Rate .$222_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$250_{\pm.090}.$$181_{\pm.039}.$$181_{\pm.039}.$$250_{\pm.034}.$
Reasoning .$409_{\pm.004}.$$369_{\pm.044}.$$453_{\pm.008}.$$539_{\pm.027}.$
Relations .$473_{\pm.028}.$$568_{\pm.048}.$$439_{\pm.033}.$$542_{\pm.027}.$
Manna (6 players, 3 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$333_{\pm.471}.$
Objective .$111_{\pm.157}.$$000_{\pm.000}.$$556_{\pm.157}.$$444_{\pm.157}.$
Reasoning .$463_{\pm.052}.$$426_{\pm.026}.$$593_{\pm.026}.$$648_{\pm.069}.$
Relations .$444_{\pm.045}.$$407_{\pm.069}.$$333_{\pm.045}.$$444_{\pm.045}.$
Overall Win Rate .$127_{\pm.059}.$$063_{\pm.022}.$$111_{\pm.045}.$$222^{*}_{\pm.081}.$
Objective .$160_{\pm.020}.$$123_{\pm.008}.$$162_{\pm.024}.$$288^{*}_{\pm.021}.$
Reasoning .$343_{\pm.006}.$$349_{\pm.012}.$$384_{\pm.002}.$$423_{\pm.009}.$
Relations .$425_{\pm.008}.$$473_{\pm.010}.$$405_{\pm.006}.$$471_{\pm.009}.$
Overall .$324_{\pm.004}.$$329_{\pm.008}.$$347_{\pm.006}.$$407^{*}_{\pm.006}.$
Table 2: Results of Agent-vs-Agent Evaluation. An “*” indicates statistical significance under the two-sample t-test with a significance level of $\alpha=0.05$, comparing it with the second-best model. More detailed data is provided in Table 8.
@@ 4.1 Experimental Setup
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models , using gpt-35-turbo-16k 0613 and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and standard deviation. In this paper, we focus on the strategy for the good camp. The murderer agent’s questioning strategy mirrors that of other agents, with an added prompt instructing it to act as if it is not the murderer while questioning others.
4.2 Evaluation Metrics
For evaluation, we adopt the following metrics:
Win Rate:
Every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) is eliminated. Murderers always vote for others instead of themselves. If the true murderer is voted out, the game is considered a victory for the identifying players.
Question Accuracy:
The accuracy of agents’ answers to three types of questions—Objective, Reasoning, and Relations—and Overall, where Overall represents the combined accuracy across all question types to provide a comprehensive evaluation of the agents’ performance. This metric is used in both Agent-vs-Agent and Human-vs-Agent evaluations to measure agents’ ability to answer inferential questions correctly.
4.3 Baselines
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games. Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games.
(1) Werewolf (Xu et al., 2023a). Werewolf is another multi-agent game, where players identify werewolves through group discussion. Questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. We adapt its pre-set game instructions and role-specific information to MMGs settings. (2) Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b). Agents think, reflect, and choose who and what to ask based on their objectives. We use the framework from previous works, only replacing the agents’ objectives with those set in MMGs. (3) ThinkThrice (Wu et al., 2024). Designed for MMGs, agents craft questions from retrieved memory and the current scenario. (4) Personal Perspective (PP). For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script. (5) Omniscient Perspective (OP). Agents do not actively participate but make their final decisions based on all agents’ scripts.
To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and Questum). The implementation details, model comparison, information on sensors, method to calculate the Overall, and prompts are provided in Appendix C.
4.4 Agent-vs-Agent Evaluation
@@
In this section, we evaluate agent performance in an Agent-vs-Agent setting to compare their decision-making processes against competitive baselines. Figure 3 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table 2). Questum  shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Notably, Questum  significantly outperforms others in Objective Questions. The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies. It represents the ideal “search” endpoint. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model. This showcases the “approximate” ability of Questum  in refining information and dynamically narrowing down the search domain to achieve the target objective. Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
@@
The results demonstrate that Questum  outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability. The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals. This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
We also evaluate the performance of our framework on English scripts and open-source LLMs in Appendix D. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
4.5 Agent-vs-Human Evaluation
In the Agent-vs-Human Evaluation, we identified significant limitations of traditional evaluation methods used in agent for games, which were primarily designed for Agent-vs-Agent Evaluation and focused solely on achieving victory. While effective in assessing agents’ strategic and reasoning abilities, these methods fall short when considering human players as the end-users of the system. For instance, previous approaches often resulted in repetitive and meaningless conversations, which human players found unengaging and boring. To illustrate these limitations, we provide specific examples in Appendix E, drawn from dialogues that occurred during games involving human participants.
To address this, we extended our evaluation by incorporating a player-centric perspective. In addition to the performance metrics used in Agent-vs-Agent Evaluation, we distributed a survey to human participants to capture their gameplay experience and satisfaction levels. This allowed us to assess Questum’s suitability not only as a strategic agent but also as a companion for human players in interactive settings. The metrics include:
Story Advancement: Measures the agent’s effectiveness in gathering relevant information and advancing the game’s plot in a meaningful way. Higher scores indicate that the interactions were productive and kept the game engaging.
Question Quality: Evaluates the relevance and depth of the questions posed by the agent. This metric ensures that the questions contribute to the narrative and help uncover critical information, avoiding redundancy or triviality.
Response Quality: Assesses the accuracy and informativeness of the agent’s responses. Higher scores indicate that the agent provided clear and valuable information to the human players.
Response Speed: Represents how quickly the agent interacts, ensuring timely response for a smooth and engaging gameplay experience.
Role Immersion: Assesses how well the agent embodies its character’s personality, emotions, and background to create a believable experience.
Evaluation Werewolf O-CoT ThinkThrice Questum
Win Rate .$111_{\pm.157}.$$333_{\pm.132}.$$333_{\pm.048}.$$667_{\pm.085}.$
Objective .$222_{\pm.091}.$$333_{\pm.045}.$$296_{\pm.052}.$$556_{\pm.091}.$
Reasoning .$678_{\pm.028}.$$644_{\pm.028}.$$684_{\pm.008}.$$723_{\pm.035}.$
Relations .$745_{\pm.028}.$$797_{\pm.040}.$$745_{\pm.028}.$$843_{\pm.042}.$
Overall .$608_{\pm.006}.$$619_{\pm.023}.$$625_{\pm.009}.$$717_{\pm.038}.$
Table 3: Performance comparison of different algorithms against human players.
Figure 4: Results of Agent-vs-Human Evaluation using Human-Centric metrics. The detailed data can be found in Table 6.
All scores are based on discrete 1-5 scales, where higher scores are better. The specific scoring criteria can be found in Table 7. Details about the setups (e.g., recruitment of players, gameplay conditions, and tasks) can be found in Appendix B.5.
As shown in Table 3 (results against human players) and Figure 4 (Human-Centric metrics), Questum  not only surpassed the baseline models in strategic and reasoning-oriented evaluation but also demonstrated superior performance in human-centric dimensions. These results indicate that Questum  effectively balances tactical prowess with narrative engagement, delivering more satisfying gameplay experiences for human participants.
4.6 Efficiency and Cost Analysis
@@
As shown in Figure 5, we delve into the efficiency and cost analysis across various methodologies for agent interaction in MMGs. The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense. [^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Figure 5: Comparison of time (hours) and costs (\$) for calling OpenAI API across multi-agent algorithms in MMG settings.
@@
We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps. In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives. Questum  stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, Questum  minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with Questum  consistently registering lower costs in comparison to its counterparts during the Gameplay phase. The trend in time efficiency mirrors that of the monetary cost, with similar patterns observed across the methodologies.